OpenStudy (anonymous):

The midpoint of is (5, –11). The coordinates of one end UV point are U(3, 5). Find the coordinates of endpoint V

OpenStudy (anonymous):

the midpoint is the average of two point. \[(\frac{x_2-x_1}{2},\frac{y_2-y_1}{2})\]substitute point U and the midpoint to find the other point

OpenStudy (amistre64):

adding, not subtracting

OpenStudy (amistre64):

if anything, the mispoint formula can be reworked as: \[x_1+x_2=2~x_m\]\[x_1=2~x_m-x_2\]

OpenStudy (anonymous):

OH ya sorry my bad

OpenStudy (anonymous):

Since the midpoint is (5, –11), it implies the total length is 2(5, -11) = (10, -22) But U(3, 5). This implies V(x, y) = (10, -22) + (3, 5) = (13, -17) Therefore the coordinate of V is (13, -17)